Stock Dividends and Stock Splits

Whitacre Company's Stockholders' Equity section of the balance sheet on December 31, 2016, was as follows:

Common stock, $10 par value, 60,000 shares $600,000
   issued and outstanding
Additional paid-in capital 480,000
Retained earnings 1,240,000
   Total stockholders’ equity $2,320,000

On May 1, 2017, Whitacre declared and issued a 15% stock dividend, when the stock was selling for $20 per share. Then on November 1, it declared and issued a 2-for-1 stock split.

Required:

1. How many shares of stock are outstanding at year-end?
shares

2. What is the par value per share of these shares?
$ per share

3. Develop the Stockholders' Equity category of Whitacre's balance sheet as of December 31, 2017.

Whitacre Company
Partial Balance Sheet - Stockholders' Equity Section
December 31, 2017
$
Total stockholders' equity $
Answer #1

1.) Stock outstanding = 60000*1.15*2 = 138000 Shares

2) Par value per share = 10/2 = 5 per share

3)

Whitacre Company
Partial Balance Sheet - Stockholders' Equity Section
December 31, 2017
Common Stock $690000
Additional paid in Capital 570000
Retained earnings 1060000
Total stockholders' equity $2320000